OpenStudy (anonymous):

5. Given the function f(x) = log4(x + 8), find the value of f^−1(2). A. f−1(2) = 4 B. f−1(2) = 8 C. f−1(2) = 10 D.f−1(2) = 17

OpenStudy (anonymous):

c

OpenStudy (anonymous):

how did you get the answer @gabriel54321 ?? I don't just want to cheat I wan't to learn.. @phi

OpenStudy (phi):

\[ f^{-1}(2) \] means find the value of f_inverse when it is given 2 as its input. they don't give you f_inverse, they give you f(x).

OpenStudy (anonymous):

so \[f^-1 (2) = \log_{4}(x+8 \]

OpenStudy (phi):

I just memorize that \( f^{-1}(2)\) means we can replace f(x) with 2 and solve for x using \[ f(x) = \log_4(x + 8) \\ 2 = \log_4(x + 8)\]

OpenStudy (anonymous):

so it would equal 8

OpenStudy (phi):

we use the idea that the "input" to f_inverse is the same as the output of f() \[ f^{-1}(a) = b \leftrightarrow f(b)= a \]

OpenStudy (anonymous):

f^-1 (2) = 8

OpenStudy (phi):

to solve we do this \[ 2 = \log_4(x+8) \] make each side the exponent to the base 4 \[ 4^2 = 4^{\log_4(x+8)} \\16= x+8 \\ 8=x \]

OpenStudy (anonymous):

thank you